OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)?

OpenStudy (anonymous):

come on, the first step is to find the slope

OpenStudy (anonymous):

first find slope\[m=\frac{5-2}{1-0}=3\]then sub into \[y-y_{1}=m(x-x _{1})\]where either of the two given points may be subbed in\[y-2=3(x-0)\]solve for y\[y=3x+2\]

OpenStudy (anonymous):

I will go just a bit further, but there is no way I am giving you the answer to this one. Sorry.. "The point-slope for with slope m in the Cartesian plane is given by y-y_1 = m(x-x_1)."

OpenStudy (anonymous):

weel someone did, just remember m=\[x_{1}-x _{0} \over y_{1}-y _{0}\]